# The James Webb Interferometer: Space-based interferometric detections of PDS 70 b and c at 4.8 m

### v1

@@ 4.1 Derived Planet Parameters
Figure 2 shows the two planet plus disk model generated from the median parameters from the uniform planet location model posterior. The 1, 2, 3, 4 and 5 contours from the marginalized posterior of the planet locations are shown by the white contours. A comparison of the model squared visibilities and closure phases to the data is given in Figure 3, showing that the model provides an excellent fit. Furthermore, the measured positions of both PDS 70 b and c are consistent with VLTI/GRAVITY predictions from Wang et al. (2021b), which are shown by the dashed grey circles in Figure 2.
Table 1 summarizes the derived planet parameters, from both of the models. The disk model parameters are listed in Table 2 in Appendix A. We find that both planets are found at a high signal-to-noise (SNR) ratio, in locations that are consistent between the uniform prior and tightly constrained Gaussian prior models. We also find that the measured contrasts of the planets are consistent for the uniform prior and GRAVITY prior models, with nearly identical contrasts derived for PDS 70 b, and a small difference for PDS 70 c, at the level.
Figure 4 shows a corner plot of the planet parameters from the uniform prior model, along with a correlation matrix of all of the model parameters. We note that there are correlations between the parameters of both planets. Notably, there is a moderate anti-correlation between the separation and the contrast of PDS 70 c, which may explain the slight discrepancy between the derived contrast of PDS 70 c between the uniform prior model and the GRAVITY prior model. From the correlation matrix, there are no strong correlations between the disk parameters and the planet contrasts, indicating that the disk model has a minimal effect on the measured contrasts.
@@ 4.2 Residual Signal and Contrast Limits
We compute an SNR map, shown in Figure 5, made by dividing the best fit point source binary contrast to the residual closure phase signal by its associated uncertainty fitting, using our new python package drpangloss444https://github.com/benjaminpope/drpangloss. drpangloss provides similar capabilities to the packages CANDID (Gallenne et al., 2015) and fouriever (Kammerer et al., 2023) but is written using JAX to provide just-in-time (JIT) compilation, GPU compatibility and auto-differentiation.
@@
Because closure phases are non-linear, we calculate the complex visibilities of the uniform prior model, using the median of the marginalized posterior for all parameters, and add this to our binary model visibilities before calculating the model closure phases. Each cell in the grid shows the best fit contrast divided by the uncertainty in the contrast calculated using the Laplace approximation (Tierney & Kadane, 1986), assuming the uncertainties are Gaussian. We clearly see that there is residual emission at an SNR of 4, to the south-west of PDS 70 A and very close to the inner working angle of 94 mas, that is not fully captured by our two planet plus disk model. To assess the nature of the signal that we observe, we use nested sampling to fit a star plus planet model to the residual closure phases and find a contrast of magnitudes, a separation of mas and a position angle of degrees. The position angle of the signal is well constrained, however, there is a strong anti-correlation between the separation and contrast, which is expected for a high contrast source near the diffraction limit. Interestingly, the position angle of the observed signal is not consistent with forward scattering from an inner disk that has the same geometry as the outer disk. However, due to the unconstrained nature of the separation of the emission it is unclear whether the signal is due to compact emission or an inner disk feature.
@@
To assess the limits that we are able to place on the presence of additional planets in the system, we calculate a 5 contrast curve, from the residual closure phase signal of the uniform prior model. We use the approach described above to fit to the residual closure phase signal, and calculate the contrast curve using the Absil method (Absil et al., 2011; Gallenne et al., 2015), as is implemented in drpangloss. We note that the Absil method is a standard approach to measuring contrast limits for interferometric data, however, it is based on a model comparison between a single star model and a binary model, and calculates the contrast at which the binary model deviates by a certain probability from the single star model (Kammerer et al., 2023). In this work, we wish to calculate an upper limit on the flux, at each location, in our field-of-view. To measure Bayesian upper limits, we adapt the method presented by Ruffio et al. (2018) to interferometric data. We first, calculate a grid of likelihoods for a binary model at fixed R.A. and Dec. locations for a range of contrasts. We next, use the maximum likelihood point in each grid cell to initialize gradient descent, using the BFGS algorithm (Broyden, 1970; Fletcher, 1970; Goldfarb, 1970; Shanno, 1970), to find the contrast that maximizes the likelihood. We use the calculated contrast in each cell, along with the Laplace approximation to calculate the uncertainty in the calculated contrast at each R.A. and Dec. point. Finally, we plug the calculated contrasts, uncertainties and a 99.999971% cutoff probability into Equation 8 in Ruffio et al. (2018), to calculate the 5 upper limit at each point in our field of view, which azimuthally average over to generate a contrast curve. We henceforth refer to this method as the Ruffio method.
@@
The derived 5 contrast/upper limit curves, and their uncertainties, calculated using the Absil and Ruffio methods, are shown in Figure 6. We reach a contrast upper limit of 7.5 mag, which is reasonably consistent with the expected photon noise limit of 7.8, given by Equation 10 by Sivaramakrishnan et al. (2023). The contours show the 1, 2 and 3 confidence intervals on the contrast and the separation for PDS 70 b and c calculated from the uniform prior model posterior.
@@
We also calculate a 5 upper limit on any flux at the reported position of the planet candidate PDS 70 d from Christiaens et al. (2024) (sep. = 115.2 mas, P.A. = 291.8∘), as our observations are only separated by 12 days. We first calculate the contrast posterior distribution using dynesty, with uniform priors on the contrast, from -1 to 1. We find a mean contrast of . To convert this measured contrast to an upper limit, we use Equation 8 by Ruffio et al. (2018), with the measured mean, standard deviation and a cutoff probability of 99.999971%, which is equivalent to a 5 upper limit, with a positivity prior on the contrast. From this, we calculate a 5 upper limit of 7.56 mag, corresponding to a flux of W m-2 m-1. We also calculate the 5 upper limit using the Laplace approximation and find an upper limit of 7.57 mag, consistent with the nested sampling result within 1%.
@@ 4.3 SED Fitting
We fit the the available data points of both protoplanets including the new NIRISS/AMI F480M measurement. For planet b, we use the SPHERE/IFS spectrum and SPHERE/IRDIS photometries from Müller et al. (2018), the VLTI/GRAVITY spectrum from Wang et al. (2021b), the VLT/NaCo photometries at m from Stolker et al. (2020a), and the F187N and F480M NIRCam measurements from Christiaens et al. (2024). For planet c, we consider the SPHERE/IRDIS and VLT/NaCo photometries from Stolker et al. (2020a), the SPHERE/IFS and VLTI/GRAVITY from Wang et al. (2021b), and the F187N and F480M NIRCam measurements from Christiaens et al. (2024).
@@
For planet b, the best fit without a CPD component is reported by the green solid line in Figure 7. The spectrum reproduces well the IFS spectrum, the SPHERE band photometries, and the F187N measurement obtained with NIRCam. However, the first half of the GRAVITY spectrum is underestimated, while the second appears to be overestimated. Finally, at m, almost none of the datapoints appear to agree within with the best-fit model, with the exception of the NIRCam F480M photometry, which has a very large uncertainty. When including the CPD contribution, the second half of the GRAVIITY spectrum is better reproduced, while at longer wavelengths the predicted flux is closer to the measurements (see orange dashed line in Figure 7). In terms of planet model parameters, the presence of MIR excess emission in the modeling suggests a hotter and smaller planet. The reduced chi square is 0.9 for the extincted Drift-PHOENIX model alone and 0.53 for the extincted Drift-PHOENIX+blackbody model. Additionally, the Bayes factor comparing the blackbody model to the no blackbody model is . Hence, the presence of the blackbody provides a better description of the observations.
The results for planet c are displayed in Figure 8. As reported by Christiaens et al. (2024), the F187N shows an excess emission, possibly due to Pa emission. The blue region of the GRAVITY spectrum does not agree with the fit, and the measurement is off by almost . Also, our NIRISS/AMI F480M measurement is clearly at odds with respect to the green solid line model fit, which does not include a CPD component. The spectrum of the best model once including the CPD presents a lower flux, consistent with the NaCo measurement. Furthermore, the CPD contribution passes right in between the F480M measurements from NIRCam and NIRISS. The reduced chi squared for these fits are 0.55 and 0.52 respectively, and the Bayes factor in favour of the model including the blackbody is 5.6, suggesting moderate support for including the blackbody component. Furthermore, assuming that the NIRISS/AMI long-wavelength measurement is less biased, weighting it more strongly would obtain a warmer/larger CPD contribution.
@@ 5.1 PDS b and c F480M Photometry
For PDS 70 b, and the uniform prior model, we find a contrast of ( mags). Similarly, for PDS 70 c, and the uniform prior model, we find a contrast of ( mags). While our result for PDS 70 b is consistent with the NIRCam F480M results of published by Christiaens et al. (2024), and reasonably consistent with the contrast of 6.15 0.27 mags published by Stolker et al. (2020a), the PDS 70 c value is inconsistent with the NIRCam F480M results of , at roughly the 3 level.
@@
There are several possible reasons for the PDS 70 c discrepancy between NIRCam and NIRISS/AMI. It is possible that the contrasts published by Christiaens et al. are systematically underestimated due to self/over subtraction, as a result of the small roll angle of 5 degrees between their two observations. Another possibility for an underestimated brightness of PDS 70 c could be due to over subtraction of the spatially coincident disk. Christiaens et al. use a radiative transfer model based on models published by Portilla-Revelo et al. (2022) and Portilla-Revelo et al. (2023), that were originally fit to ALMA data and 1.25 m polarized VLT/SPHERE data of PDS 70. These models were iterated on by fitting for the minimum grain size and settling parameters to less than half of the forward scattering side of the disk, in F480M. This tightly constrained fitting approach could lead to biased results that do not adequately reproduce the disk emission at the location of PDS 70 c. Specifically, an incorrect estimation of the scattering phase function will significantly impact the PDS 70 c measured contrast.
@@
Separating out the disk emission is less of an issue for the interferometric data set despite the simple disk model employed, using a symmetric, optically thick disk component plus the asymmetric Henyey-Greenstein scattering phase function. The signals of the disk and the planet in Fourier space are well separated and thus we can disentangle the asymmetric contributions from the forward scattering peak of the disk, and PDS 70 c, offset by 20 degrees, due to the majority of our likelihood being from the closure phases (the squared visibilities are nearly flat at a signal level of 0.90, Figure 3). Furthermore, our joint modelling of the planet and disk parameters allows us to measure the correlations between the parameters of both planets and the disk, which are shown by the correlation matrix in the top right of Figure 4. Notably for the contrast of PDS 70 c, there is a moderate correlation with g, the scattering parameter (more forward scattering leads to less emission), a strong correlation with the separation of PDS 70 b (larger separation leads to less emission), a moderate anti-correlation with , the disk inner-edge truncation parameter and a moderate anti-correlation with its own separation from the star (plus unresolved inner disk emission).
@@
We note that these correlations, between the contrast of PDS 70 c and the disk model parameters, are small compared to the correlation between the contrast of PDS 70 c and the separation of PDS 70 b, which is evident in Figure 4. This means a larger separation of PDS 70 b would correspond to a fainter contrast of PDS 70 c. To asses the impact of this correlation, we calculate the contrast of PDS 70 c using only posterior samples within 3 of the predicted location of PDS 70 c from the orbit fitting presented in Wang et al. (2021b). We find a contrast of 6.53 0.12 mags, consistent with our value of 6.48 0.13 mags, calculated using the full posterior. As an additional test to explore the impact of the scattering phase function profile, we briefly explored other simple parametric scattering phase functions. We tested an azimuthal Gaussian profile and the sum of an azimuthal Gaussian profile and the Henyey-Greenstein scattering phase function. We repeated the model fitting procedure described in Section 3, using uniform priors on the planet locations, and found contrasts for PDS 70 b and c that were nearly identical to the results reported in Table 1 (and Table 3, in Appendix A) for both models.
@@
As another test, we briefly explored including an inner disk component in our model, with its geometry fixed to that of the outer disk. We tested symmetric models as well as models that included skewed emission in the forward scattering direction. We found that symmetric models did not noticeably improve the fit. For the skewed models, we found that the fit did somewhat improve, for a large, bright, extended inner disk model, however, we saw significant correlations between the inner disk and the planet location parameters. Such a bright and large inner disk is inconsistent with the size and mass of the inner disk that has been inferred from SED fitting (e.g., Gaidos et al., 2024). Additionally, the planet contrasts that we found from these tests were all consistent with the planets plus outer disk only model, well within of the contrasts that we report in Table 3. Given these points, as well as the fact that the inner disk is marginally resolved, if at all, and that it possibly has a complicated morphology (Mesa et al., 2019; Casassus & Cárcamo, 2022), not well captured by a geometrical model, we decided not to further explore an inner disk model component.

### v2

@@ 4.1 Derived Planet Parameters
Figure 3 shows the two planet plus disk model generated from the median parameters from the power law scattering model (Equation 8) model posterior. We focus on the power law model in particular, because it had the highest Bayesian evidence out of the three models we explored (Table 2 in the Appendix). The 1, 2 and 3 contours from the marginalized posterior of the planet locations are shown by the white contours. A comparison of the model log kernel amplitudes and kernel phases to the data is given in Figure 4, showing that the model provides an excellent fit. Furthermore, the measured positions of both PDS 70 b and c are consistent with VLTI/GRAVITY predictions from Wang et al. (2021b), which are shown by the dashed grey circles in Figure 3.
Table 1 summarizes the derived planet parameters, calculated from the Bayesian average of all three models. The disk model parameters are listed in Table 2 in Appendix A. We find that both planets are found at a high signal-to-noise ratio (SNR), 14.7 for PDS 70 b and 7.0 for PDS 70 c (where we define the SNR as the contrast divided by the uncertainty in the contrast marginalized over all of the model parameters), in locations that are consistent between the three models (planet parameters for the individual models are shown in Table 3 in the Appendix). We also find that the measured contrasts of the planets are consistent between the three models.
Figure 5 shows a corner plot of the planet parameters for each of the three models, as well as the Bayesian average of the three models. We note that there are correlations between the parameters of both planets. Notably, there is a moderate anti-correlation between the separation and the contrast of PDS 70 c. From the correlation matrices of the model parameters, shown in Figure 6, there are no strong correlations between the disk parameters and the planet contrasts, indicating that the disk model has a minimal effect on the measured contrasts.
| PDS 70 b | PDS 70 c | |
| Separation (mas) | ||
| Position Angle (∘) | ||
| Contrast (mag) | ||
| Flux1 (Jy) |
Notes: We report the median along with the 16th and 84th percentiles of the Bayesian average of the marginalized posteriors over the three disk (Equations 6, 7, 8) plus plus star planets models we explored.
1 We convert the planet contrasts to flux using our measured star plus unresolved inner disk flux of mJy.
@@ 4.2 Residual Signal and Contrast Limits
We compute an SNR map, shown in Figure 7, made by dividing the best fit point source binary contrast to the residual kernel phase signal by its associated uncertainty fitting.
@@
Because closure phases are non-linear, we calculate the complex visibilities of the power law model, using the median of the marginalized posterior for all parameters of the power law model, and add this to our binary model visibilities before calculating the model closure phases and then kernel phases. Each cell in the grid shows the best fit contrast divided by the uncertainty in the contrast calculated using the Laplace approximation (Tierney & Kadane, 1986), assuming the uncertainties are Gaussian. We clearly see that there is residual emission at an SNR of 4, to the south west of PDS 70 A and very close to the inner working angle of 94 mas, that is not fully captured by our two planet plus disk model. To assess the nature of the signal that we observe, we use nested sampling to fit a star plus planet model to the residual kernel phases and find a contrast of magnitudes, a separation of mas and a position angle of degrees. The position angle of the signal is reasonably well constrained, however, there is a strong correlation between the separation and contrast, which is expected for a high contrast source near the diffraction limit. Interestingly, the position angle of the observed signal is not consistent with forward scattering from an inner disk that has the same geometry as the outer disk. However, due to the unconstrained nature of the separation of the emission it is unclear whether the signal is due to compact emission or an inner disk feature.
@@
To assess the limits that we are able to place on the presence of additional planets in the system, we calculate a 5 contrast curve, from the residual kernel phase signal of the power law model. We use the approach described above to fit to the residual kernel phase signal, and calculate Bayesian upper limits. To do this, we adapt the method presented by Ruffio et al. (2018) to interferometric data. We first calculate a grid of likelihoods for a binary model at fixed R.A. and Dec. locations for a range of contrasts. We next use the maximum likelihood point in each grid cell to initialize gradient descent, using the BFGS algorithm (Broyden, 1970; Fletcher, 1970; Goldfarb, 1970; Shanno, 1970), to find the contrast that maximizes the likelihood. We use the calculated contrast in each cell, along with the Laplace approximation to calculate the uncertainty in the calculated contrast at each R.A. and Dec. point. Finally, combining the calculated contrasts, uncertainties and a 99.999971% (equivalent to a ) cutoff probability and using Equation 8 in Ruffio et al. (2018), we calculate the 5 upper limit at each point within a radius of 400 mas, which we azimuthally average over to generate a contrast curve. We henceforth refer to this method as the Ruffio method.
@@
The derived 5 contrast/upper limit curves, along with the 1 and 3 curves, calculated using the Ruffio method, are shown in Figure 8. Beyond 110 mas, we reach a 5 contrast upper limit of 7 magnitudes, and a 3 limit of 7.6-7.8 magnitudes, which is consistent with the expected photon noise limit of 7.8, given by Equation 10 by Sivaramakrishnan et al. (2023). The contours show the 1, 2 and 3 confidence intervals on the contrast and the separation for PDS 70 b and c calculated from the Bayesian average of the three model posteriors.
@@
We also calculate a 5 upper limit on any flux at the reported position of the planet candidate PDS 70 d from Christiaens et al. (2024) (sep. = 115.2 mas, P.A. = 291.8∘), as our observations are only separated by 12 days. We first calculate the contrast posterior distribution using dynesty, with uniform priors on the contrast, from -1 to 1. We find a mean contrast of . To convert this measured contrast to an upper limit, we use Equation 8 by Ruffio et al. (2018), with the measured mean, standard deviation and a cutoff probability of 99.999971%, which is equivalent to a 5 upper limit, with a positivity prior on the contrast. From this, we calculate a 5 upper limit of 7.04 mag, corresponding to a flux of 208 10 Jy. We also calculate the 5 upper limit using the Laplace approximation, again retrieving an upper limit of 7.04 mag, in line with our assumption of a Gaussian posterior.
@@ 4.3 SED Fitting
We fit the available data points of both protoplanets including the new NIRISS/AMI F480M measurement. For planet b, we use the SPHERE/IFS spectrum and SPHERE/IRDIS photometry from Müller et al. (2018), the VLTI/GRAVITY spectrum from Wang et al. (2021b), the VLT/NaCo photometry at m from Stolker et al. (2020), and the F187N and F480M NIRCam measurements from Christiaens et al. (2024). For planet c, we consider the SPHERE/IRDIS and VLT/NaCo photometry from Stolker et al. (2020), the SPHERE/IFS and VLTI/GRAVITY from Wang et al. (2021b), and the F187N and F480M NIRCam measurements from Christiaens et al. (2024).
@@
For planet b, the best fit without a CPD component is reported by the green solid line in Figure 9. The spectrum reproduces well the IFS spectrum, the SPHERE band photometry, and the F187N measurement obtained with NIRCam. However, the first half of the GRAVITY spectrum is underestimated, while the second appears to be overestimated. Finally, at m, almost none of the data points appear to agree within with the best-fit model, with the exception of the NIRCam F480M photometry, which has a very large uncertainty. When including the CPD contribution, the second half of the GRAVIITY spectrum is better reproduced, while at longer wavelengths the predicted flux is closer to the measurements (see orange dashed line in Figure 9). In terms of planet model parameters, the presence of MIR excess emission in the modeling suggests a hotter and smaller planet. The Bayes factor comparing the blackbody model to the no blackbody model is . Hence, the presence of the additional blackbody emission provides a better description of the observations.
The results for planet c are displayed in Figure 10. As reported by Christiaens et al. (2024), the F187N shows an excess emission, possibly due to Pa emission. The blue region of the GRAVITY spectrum does not agree with the fit, and the measurement is off by almost . Also, our NIRISS/AMI F480M measurement is clearly at odds with respect to the green solid line model fit, which does not include a CPD component. The spectrum of the best model once including the CPD presents a lower flux, consistent with the NaCo measurement. Furthermore, the CPD contribution passes right in between the F480M measurements from NIRCam and NIRISS. The Bayes factor of the model including the blackbody compared with the blackbody-free model is 0.88, indicating that the model without the additional blackbody component is preferred. However, we note that our measurement of PDS 70 c hints that there is in fact excess emission that can’t be explained by the atmosphere only model, and the fit is being constrained by the fainter measurement of PDS 70 c made by Christiaens et al. (2024). Follow-up observations at 4.8 m with NIRISS and NIRCam, and beyond 5 m with MIRI, could be used to confirm the presence of warm CPD emission and constrain the CPD properties.
@@ 5.1 PDS b and c F480M Photometry
For PDS 70 b, we find a F480M flux of Jy. Similarly, for PDS 70 c, we find a flux of Jy. While our result for PDS 70 b is consistent with the NIRCam F480M results of Jy published by Christiaens et al. (2024), the PDS 70 c value is significantly inconsistent with the NIRCam F480M results of Jy.
@@
There are several possible reasons for the PDS 70 c discrepancy between NIRCam and NIRISS/AMI. It is possible that the fluxes published by Christiaens et al. (2024) are systematically underestimated due to self/over subtraction, as a result of the small roll angle of 5 degrees between their two observations. Another possibility for an underestimated brightness of PDS 70 c could be due to over subtraction of the spatially coincident disk. Christiaens et al. (2024) use a radiative transfer model based on models published by Portilla-Revelo et al. (2022) and Portilla-Revelo et al. (2023), that were originally fit to ALMA data and 1.25 m polarized VLT/SPHERE data of PDS 70. These models were iterated on by fitting for the minimum grain size and settling parameters to less than half of the forward scattering side of the disk, in F480M. This tightly constrained fitting approach could lead to biased results that do not adequately reproduce the disk emission at the location of PDS 70 c. Specifically, an incorrect estimation of the scattering phase function will significantly impact the PDS 70 c measured contrast. It is also possible that the flux of PDS 70 presented in this work is over estimated, due to emission from the spiral structure that was tentatively detected in Christiaens et al. (2024), that is not included in our geometrical model. However, any bias from not including this component in our model should be minimal, due to its extended nature. A final explanation for the significant difference between our flux measurements of PDS 70 c is that the flux of PDS 70 c itself is variable at 4.8 m, due to non-steady accretion from the circumplanetary material.
@@
Separating out the disk emission is less of an issue for the interferometric data set despite the simple disk models employed, using a symmetric, optically thick disk component plus the asymmetric Gaussian, power law and Henyey-Greenstein azimuthal brightness profiles. The signals of the disk and the planet in Fourier space are well separated and thus we can disentangle the asymmetric contributions from the forward scattering peak of the disk, and PDS 70 c, offset by 20 degrees.Furthermore, our joint modelling of the planet and disk parameters allows us to measure the correlations between the parameters of both planets and the disk, which are shown by the correlation matrices in Figure 6. Notably for the contrast of PDS 70 c, there are moderate correlations with the model parameters that control the azimuthal brightness profile of the disk, , g and (more forward scattering leads to less emission), a strong correlation with the separation of PDS 70 b (larger separation leads to less emission), a moderate anti-correlation with , the disk inner-edge truncation parameter, and a moderate anti-correlation with its own separation from the star (plus unresolved inner disk emission).
@@
We note that these correlations, between the contrast of PDS 70 c and the disk model parameters, are small compared to the correlation between the contrast of PDS 70 c and the separation of PDS 70 b, which is evident in Figure 6. This means a larger separation of PDS 70 b would correspond to a fainter contrast of PDS 70 c. To asses the impact of this correlation, we calculate the contrast of PDS 70 c using only posterior samples within 3 of the predicted location of PDS 70 c from the orbit fitting presented in Wang et al. (2021b). We find a contrast of mags, consistent with our value of mags, calculated using the full posterior.
@@
As another test, we briefly explored including an inner disk component in our models, with its geometry fixed to that of the outer disk. We tested symmetric models as well as models that included skewed emission in the forward scattering direction. We found that symmetric models did not noticeably improve the fit. For the skewed models, we found that the fit did somewhat improve, for a large, bright, extended inner disk model, however, we saw significant correlations between the inner disk and the planet location parameters. Such a bright and large inner disk is inconsistent with the size and mass of the inner disk that has been inferred from SED fitting (e.g., Gaidos et al., 2024). Additionally, the planet contrasts that we found from these tests were all consistent with the planets plus outer disk only model, well within of the contrasts that we report in Table 3. Given these points, as well as the fact that the inner disk is marginally resolved, if at all, and that it possibly has a complicated morphology (Mesa et al., 2019; Casassus & Cárcamo, 2022), not well captured by a geometrical model, we decided not to further explore an inner disk model component.